Reports 1-1 of 1 Clear search Modify search
AdV-TCS (CO2 laser projector)
bersanetti, nardecchia, zaza - 22:56 Tuesday 26 August 2025 (67571) Print this report
Lock acquisition recovery after NI CO2 bench intervention

After the activity on the NI CO2 bench, the ITF successfully relocked at CARM NULL 1F on the first attempt at 16.58 UTC.
However, we observed that the power on B1p was too high (0.03 mW) to proceed with the lock acquisition, so we started tuning the NI DAS to reduce it.

Step 1 (17:38 UTC): Decreased outer ring power

  • NI CO2 PICK OFF: 0.55 W → 0.52 W
    Positive effect on B1p

Step 2 (18:26 UTC): decreased outer ring power

  • NI CO2 PICK OFF: 0.52 W → 0.49
    ​​​​​​Negative effect on B1p

Step 3 (19:04 UTC): Increased outer ring power

  • NI CO2 PICK OFF: 0.49 W → 0.51  

Step 4 (19:20 UTC): Increased inner ring power

NI CO2 PICK OFF: 0.51 W → 0.522
​​
Once the B1p power reached 0.02 mW, we resumed the lock acquisition.
The ITF relocked in LN3 at 20:36 UTC.
Lock acquisition procedure is still to be verified.

Images attached to this report
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×